OpenStudy (anonymous):

4/x = x/16 can someone help please?

OpenStudy (anonymous):

are we finding x?

OpenStudy (paki):

cross multiply... x^2= 64 or x= 8

OpenStudy (anonymous):

sqrtroot (16*4)=x

OpenStudy (anonymous):

yes! @noneyabusiness

OpenStudy (anonymous):

x = 64/x

OpenStudy (anonymous):

you would do what @paki said

OpenStudy (anonymous):

x=-8 x=8

OpenStudy (anonymous):

@mdh227 Did that help?

OpenStudy (paki):

"x" will be only +8.... not -8...

OpenStudy (anonymous):

yea yea yea!!! @Night-Watcher

OpenStudy (anonymous):

Use this site, it will be most helpful: https://www.wolframalpha.com/